OpenStudy (anonymous):

Take the region between the curves y = x and y = x^2 and revolve it around the line y = 2. Show work. the answer is (8*pi)/15

OpenStudy (anonymous):

lets find the limits of integration: x=x^2 x^2-x=0 x(x-1)=0 which means x=0 and x=1

OpenStudy (anonymous):

yep got that

OpenStudy (anonymous):

so the graph looks something like this: